# Hidden ferromagnetism of centrosymmetric antiferromagnets

### v1

@@ VI.3 $\rm RuO_{2}$}-type systems
The largest contributions to $\Omega^{y}_{\boldsymbol{k}}$ arise from the $(2\pi,0,0)\to(0,0,0)\to(0,0,2\pi)$ directions in the $k_{y}=0$ plane, which are additionally modulated along $k_{y}$, as described by Eq. (44). Moreover, $\Omega^{y}_{\boldsymbol{k}}$ has nodal lines along $k_{z}=\pm k_{x}$. Therefore, if the bands along $(0,0,0)\to(2\pi,0,0)$ and $(0,0,0)\to(0,0,2\pi)$ were equally populated, $\sigma_{xy}$ would vanish. That is why we need the “tetragonal strain” $\delta t_{2}$ [26], which in the present case is nothing but the tetragonal distortion $c/a<1$ of the bct lattice. It leads to considerably stronger dispersion along $(0,0,0)\to(0,0,2\pi)$. Then, for the $n_{\rm el}=0.5$ electrons, the “majority-spin” band is fully populated along $(0,0,0)\to(2\pi,0,0)$ and only partially populated along $(0,0,0)\to(0,0,2\pi)$, resulting in finite $\sigma_{xy}$. In the expression (38) for ${\cal M}^{y}_{\boldsymbol{k}}$, $\Omega^{y}_{\boldsymbol{k}}$ is additionally modulated by $h^{\phantom{0}}_{\boldsymbol{k}}$, thus destroying the perfect cancellation in between the directions $(0,0,0)\to(2\pi,0,0)$ and $(0,0,0)\to(0,0,2\pi)$. Unlike in the previous two examples, there is a substantial deviation in the from of the anomalous Hall conductivity and orbital magnetization in Fig. 7(f). This effect is related to the much stronger tetragonal distortion, $\delta t_{2}/t_{1}=1.76$ [51], when the term $h^{\phantom{0}}_{\boldsymbol{k}}\Omega^{y}_{\boldsymbol{k}}$ in the orbital magnetization starts to play a sizable role.
VII Summary and Outlook
@@
The time-reversal symmetry breaking in AFM substances implies certain analogies with the ferromagnetism. On the one hand, it lifts the Kramers degeneracy, splitting the AFM bands with different spin quantum numbers, like in the FM systems. This is an overt analogy, which is intensively explored in the context of altermagnetism [2, 3, 4], but relies on more or less standard picture of antiferromagnetism where there are two magnetic sublattices with opposite directions of magnetic moments.
@@
In the present work, we have argued that there is a deeper analogy, which allows us to present certain classes of centrosymmetric antiferromagnets with broken $\mathcal{T}$ as if they were ferromagnets with only one magnetic site in the unit cell.
@@
Such a presentation appears to be possible due to the hidden $\{\mathcal{S}|{\bf t}\}$ symmetry of the SO interaction in the centrosymmetric antiferromagnets, specifying the transformation to the local coordinate frame where all the spins are pointed along positive direction of $z$, i.e. like in ferromagnets. This symmetry is the consequence of the inversional and translational invariance of the system. The key point is that the inversion symmetry can be broken locally, in individual bonds. Nevertheless, in the centrosymmetric antiferromagnets, $\mathcal{I}$ remains to be one of the symmetry operations of the magnetic lattice. Therefore, if the system is subjected to the lattice distortion, in order to preserve $\mathcal{I}$, this distortion must be antiferroelectric. The antiferroelectric distortion inevitably leads to the doubling of the crystallographic cell, so that the SO interaction behaves as an antiferromagnetically ordered object and changes its sign when it is calculated around the sites of different antiferroelectric sublattices. Nevertheless, the sign change of the SO interaction parameters can be compensated by the similar sign change of the spin Pauli matrices in the local coordinate frame.
@@
Thus, the $\{\mathcal{S}|{\bf t}\}$ symmetry justifies the use of the generalized Bloch theorem, which is widely employed in calculations of incommensurate spin-spiral textures without the SO coupling [28, 49]. Using the generalized Bloch theorem, any such incommensurate texture with the propagation vector $\boldsymbol{q}$ can be described within the crystallographic cell. The AFM alignment of spins is a particular case of the spin spiral, where $\boldsymbol{q}$ is the half of the reciprocal translation. However, the generalized Bloch theorem is typically at odds with the relativistic SO interaction, which tends to lock the spins along some easy axis and severely restrict their ability to rotate on the lattice [59]. The antiferromagnetism appears to be very special case: even though the generalized Bloch theorem is not valid for an arbitrary propagation $\boldsymbol{q}$, it is valid for the AFM propagation, where the antiferroelectric lattice distortion imposes a special symmetry constrain on the SO interaction. As the result, in the local coordinate frame, the SO interaction part responsible for AHE and orbital magnetism, becomes translationally invariant on the “ferromagnetic” lattice with one cite in the unit cell. The same holds for the Néel field.
@@
The regular AFM order doubles the unit cell. In unconventional antiferromagnets with broken $\mathcal{T}$, the magnetic unit cell is expected to coincide with the crystallographic one [1]. We substantially revise this canonical point of view and show that, due to the $\{\mathcal{S}|{\bf t}\}$ symmetry, the magnetic unit cell of centrosymmetric antiferromagnets can be even smaller than the crystallographic one. The situation is highly unusual, but reflects the fundamental symmetry of these systems and naturally explains their similarity with ferromagnets. The emergence of AHE and orbital magnetization in this case becomes natural and they can be evaluated in the same way as for the regular ferromagnets, but in the local coordinate frame, specified by the symmetry operation $\{\mathcal{S}|{\bf t}\}$.
@@
We would like to emphasize that the considered effect has nothing to do with the altermagnetic splitting of bands in centrosymmetric antiferromagnets. From the viewpoint of symmetry, these are different phenomena: the $\{\mathcal{S}|{\bf t}\}$ symmetry is a consequence of the spatial inversional and translational invariance, while the band splitting occurs due to rotational $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ symmetry. The latter does not break the $\{\mathcal{S}|{\bf t}\}$ symmetry of the SO interaction itself, which operates between the sites belonging to different antiferroelectric sublattices. The main effect of the $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ symmetry on the SO interaction is to fix the signs of different components of this interaction around each magnetic site. However, $\{\mathcal{T}\mathcal{C}|{\bf t}\}$ imposes a symmetry constraint also on the hopping parameters operating within the same sublattices by transforming the hoppings in one magnetic sublattice to “rotated” hoppings in another sublattice. This can split the AFM bands and eventually break the $\{\mathcal{S}|{\bf t}\}$ symmetry.
@@
If the microscopic Hamiltonian were real, the transformations $\{\mathcal{S}|{\bf t}\}$ and $\{\mathcal{T}|{\bf t}\}$ would be identical. However, when the bonds connecting different sublattices are noncentrosymmetric, as in the antiferroelectrically distorted lattice, the Hamiltonian is complex. Then, $\{\mathcal{S}|{\bf t}\}$ is not the same as $\{\mathcal{T}|{\bf t}\}$. Therefore, the time-reversal symmetry can be broken, but the Hamiltonian remain invariant under $\{\mathcal{S}|{\bf t}\}$, making the bands degenerate. This is the fundamental reason why the centrosymmetric antiferromagnets can exhibit AHE and related phenomena. The band splitting provides an additional contribution to AHE. However, it is expected to be relatively small [26].
@@
Thus, the altermagnetic splitting of bands is certainly an interesting theoretical discovery leading to a number of practically important phenomena such as the spin-current generation in AFM substances [15, 18, 24]. However, regarding the whole spectrum of properties expected in centrosymmetric antiferromagnets with broken time-reversal symmetry, it would not be right to attribute all of them to the band splitting and try to consider all of them only from the viewpoint of this splitting. Even if the band splitting is small, as was recently observed in some potential altermagnets [60], the material can still host robust AHE and net orbital magnetization, exhibit magneto-optical rotations and other “ferromagnetic” phenomena, which are related to other, hidden, symmetries of such AFM state, being the consequence if inversional and translational invariance.

### v2

@@ VI.3 $\rm RuO_{2}$}-type systems
The largest contributions to $\Omega^{y}_{\boldsymbol{k}}$ arise from the $(2\pi,0,0)\to(0,0,0)\to(0,0,2\pi)$ directions in the $k_{y}=0$ plane, which are additionally modulated along $k_{y}$, as described by Eq. (45). Moreover, $\Omega^{y}_{\boldsymbol{k}}$ has nodal lines along $k_{z}=\pm k_{x}$. Therefore, if the bands along $(0,0,0)\to(2\pi,0,0)$ and $(0,0,0)\to(0,0,2\pi)$ were equally populated, $\sigma_{xy}$ would vanish. That is why we need the “tetragonal strain” $\delta t_{2}$ [17, 26], which is associated with the tetragonal distortion $c/a<1$ of the bct lattice. It leads to considerably stronger dispersion along $(0,0,0)\to(0,0,2\pi)$. Then, for the $n_{\rm el}=0.5$ electrons, the “majority-spin” band is fully populated along $(0,0,0)\to(2\pi,0,0)$ and only partially populated along $(0,0,0)\to(0,0,2\pi)$, resulting in finite $\sigma_{xy}$. In the expression (39) for the orbital magnetization, $\Omega^{y}_{\boldsymbol{k}}$ is additionally modulated by $h^{\phantom{0}}_{\boldsymbol{k}}$, thus destroying the perfect cancellation between the directions $(0,0,0)\to(2\pi,0,0)$ and $(0,0,0)\to(0,0,2\pi)$. Unlike in the previous two examples, there is a substantial deviation in the from of the anomalous Hall conductivity and orbital magnetization in Fig. 8(f). This effect is related to the much stronger tetragonal distortion, $\delta t_{2}/t_{1}=1.76$ [51], when the term $h^{\phantom{0}}_{\boldsymbol{k}}\Omega^{y}_{\boldsymbol{k}}$ in the orbital magnetization starts to play a sizable role.
VII Classification based on realistic electronic structure
The existence of AHE without the band splitting is a new and important finding. So far it was introduced using simple one-orbital toy model [26]. This model can be justified for a limited number of compounds, like La2CuO4, VF4, and CuF2, where the electronic structure near the Fermi level is indeed featured by a single well-separated band, which can be taken as the basis for the construction of realistic model and using for these purposes the input from first-principles electronic-structure calculations [65, 66, 67, 68]. This model is typically in the strong coupling limit meaning that the transfer integrals responsible for the band splitting, $\delta t_{\boldsymbol{R},\boldsymbol{R}}\sim\varepsilon^{\uparrow}_{\boldsymbol{k}}-\varepsilon^{\downarrow}_{\boldsymbol{k}}=\delta\varepsilon_{\boldsymbol{k}}$ are much weaker than $B$, so that the contribution to the Berry curvature associated with the band splitting is about $\delta\varepsilon_{\boldsymbol{k}}/B$ times smaller than the one coming from degenerate bands [26].
The next important questions is how the contributions related to the band splitting and spin-degeneracy can be separated in realistic electronic structure calculations. For these purposes, one can use results of the model analysis as a guidance. First, it is important to stay in the global coordinate frame: although generalized Bloch theorem is vital for clarifying fundamental aspects of the time-reversal symmetry breaking in centrosymmetric antiferromagnets, it relies on additional approximations, such as the neglect of the spin-splitting of AFM bands, which do not necessarily hold in realistic calculations. Furthermore, the generalized Bloch theorem does not provide a substantial technical advantage for calculating AFM spin structures (in comparison with incommensurate spin spirals). Second, spin should be a good quantum number (otherwise, the analysis of spin-splitting is meaningless). Therefore, one should keep only one component of the SO coupling, which is parallel to the Néel field. This is similar to what we did in the model. Two other components of the SO coupling cannot be fully eliminated as in the one-orbital case. However, if the spin canting is small, their contribution to AHE is also expected to be small [16].
Suppose that the crystal is oriented in such a way that the Néel field and SO coupling are parallel to $z$. Importantly, using this setting and only $z$ component of the SO coupling, we should be able to reproduce $z$ component of DM interactions [30, 64] and, hopefully, the Hall conductivity $\sigma_{xy}$, which is related to the sign-alternating part of these interactions [26].
Finally, let us assume that the altermagnetic bands near the Fermi level can be classified by the spin-doublet quantum number $n$ and spin indices, $\uparrow$ or $\downarrow$, describing the splitting of each such doublet. Then, the Hall conductivity,
$\sigma_{xy}=-\frac{1}{V}\sum_{n}\int_{\rm BZ}\frac{d\boldsymbol{k}}{\Omega_{\rm BZ}}\left(f^{\uparrow}_{n\boldsymbol{k}}\Omega^{\uparrow,z}_{n\boldsymbol{k}}+f^{\downarrow}_{n\boldsymbol{k}}\Omega^{\downarrow,z}_{n\boldsymbol{k}}\right),$
can be rearranged as
$\sigma_{xy}=-\frac{1}{V}\sum_{n}\int_{\rm BZ}\frac{d\boldsymbol{k}}{\Omega_{\rm BZ}}\left(f^{\rm e}_{n\boldsymbol{k}}\Omega^{{\rm o},z}_{n\boldsymbol{k}}+f^{\rm o}_{n\boldsymbol{k}}\Omega^{{\rm e},z}_{n\boldsymbol{k}}\right),$
where $\Omega^{{\rm o},z}_{n\boldsymbol{k}}=\frac{1}{2}(\Omega^{\uparrow,z}_{n\boldsymbol{k}}-\Omega^{\downarrow,z}_{n\boldsymbol{k}})$ and $f^{\rm o}_{n\boldsymbol{k}}=(f^{\uparrow}_{n\boldsymbol{k}}-f^{\downarrow}_{n\boldsymbol{k}})$ are odd in the Néel field $B$ (the interchange of $\uparrow$ and $\downarrow$), while $\Omega^{{\rm e},z}_{n\boldsymbol{k}}=\frac{1}{2}(\Omega^{\uparrow,z}_{n\boldsymbol{k}}+\Omega^{\downarrow,z}_{n\boldsymbol{k}})$ and $f^{\rm e}_{n\boldsymbol{k}}=(f^{\uparrow}_{n\boldsymbol{k}}+f^{\downarrow}_{n\boldsymbol{k}})$ are even. Obviously, total $\sigma_{xy}$ is odd in $B$. Basically, this is a generalization of results of the one-orbital model [26]. The first term can be finite even when the bands are spin-degenerate, while the second term is finite only when the $\uparrow$ and $\downarrow$ bands are split across the Fermi level, like in the FM state. Indeed, according to Eq. (29), $\Omega^{z}_{n\boldsymbol{k}}$ is odd in $B$ when the bands are spin-degenerate (in the global frame), manifesting the $\{\mathcal{S}|{\bf t}\}$ symmetry. The even term, $\Omega^{{\rm e},z}_{n\boldsymbol{k}}$, appears when the $\{\mathcal{S}|{\bf t}\}$ symmetry is broken by the altermagnetic band splitting (see Table 1).
It is instructive to briefly discuss the relative importance of $\Omega^{{\rm o},z}_{n\boldsymbol{k}}$ and $\Omega^{{\rm e},z}_{n\boldsymbol{k}}$ in some characteristic materials using the simplified expression $\Omega^{{\rm e},z}_{n\boldsymbol{k}}/\Omega^{{\rm o},z}_{n\boldsymbol{k}}\sim\delta\varepsilon^{\max}_{n}/B$ anticipated in the strong coupling limit [26], where $\varepsilon^{\max}_{n}$ is a characteristic (maximal) spin-splitting near the Fermi level (a measure of electron hoppings responsible for this splitting). Particularly, an exceptionally large $\delta\varepsilon^{\max}_{n}=0.8$ eV has been reported in MnTe on the basis of ARPES measurements [20]. However, $B$ in MnTe is also expected to be large. In the local spin density approximation, it can be estimated as $B\sim\frac{1}{2}JM$, where $J$ is a characteristic exchange coupling (typically being of the order of intraatomic Hund’s rule coupling, $J\sim 0.8$ eV in the case of Mn2+ [69]) and $M$ is the local spin magnetic moment ($M\sim 4.25\mu_{\rm B}$ in MnTe [70]). Therefore, the ratio $\delta\varepsilon^{\max}_{n}/B$ can be estimated as $0.47$. The effects of on-site Coulomb repulsion $U$ are also anticipated [70, 71]. Particularly, like in the one-orbital model, $U$ can contribute to $B$ and further decrease the ratio $\delta\varepsilon^{\max}_{n}/B$.
RuO2 is an exceptional case. According to theoretical calculations [16], $\varepsilon^{\max}_{n}$ is large ($\sim 1$ eV), while $M$ is small ($\sim 1.17\mu_{\rm B}$). Furthermore, $J$ for the $4d$ materials is also expected to be relatively small ($\sim 0.6$ eV [69]). Then, $\delta\varepsilon^{\max}_{n}/B$ can be estimated as $2.85$. This value can be somewhat decreased by considering the effects of on-site $U$, which are also anticipated in RuO2 [16]. In any case, the even-order contribution $\Omega^{{\rm e},z}_{n\boldsymbol{k}}$ is expected to be large in RuO2. However, it does not necessarily mean that the odd-order contribution $\Omega^{{\rm o},z}_{n\boldsymbol{k}}$ is negligible. Furthermore, whether $\Omega^{{\rm e},z}_{n\boldsymbol{k}}$ will dominate in the calculations of $\sigma_{xy}$ depends on how the topology of $\Omega^{{\rm e},z}_{n\boldsymbol{k}}$ matches the one of the Fermi surface and whether the areas of large spin-splitting of bands across the Fermi level will coincide with the ones where $\Omega^{{\rm e},z}_{n\boldsymbol{k}}$ is large.
VIII Summary and Outlook
@@
The time-reversal symmetry breaking in AFM substances implies certain analogies with ferromagnetism. On the one hand, it lifts Kramers degeneracy and splits the AFM bands with opposite spins. This is an overt analogy, which relies on more or less standard picture of antiferromagnetism where there are two magnetic sublattices with the moments pointing in opposite directions. Today, this analogy is intensively explored in the context of altermagnetism [2, 3, 4].
@@
In the present work, we have argued that there is a deeper analogy, which allows us to present certain classes of antiferromagnets with broken $\mathcal{T}$ as if they were ferromagnets with only one magnetic site per cell.
@@
Such a presentation appears to be possible due to the hidden $\{\mathcal{S}|{\bf t}\}$ symmetry of SO interaction in centrosymmetric antiferromagnets, specifying the transformation to the local coordinate frame where all the spins are pointed along the positive direction of $z$, i.e. like in ferromagnets. This symmetry is the consequence of the inversional invariance of the system. The key point is that the inversion symmetry can be broken locally, in individual bonds. Nevertheless, in centrosymmetric antiferromagnets, some inversion centers continue to exist, so that $\mathcal{I}$ transforms the AFM lattice to itself. If the system is subjected to the lattice distortion, in order to preserve $\mathcal{I}$, this distortion must be antiferroelectric. In the other words, we are dealing with the antiferroelectric antiferromagnetism.
The antiferroelectric distortion inevitably leads to the doubling of the crystallographic cell, where the SO interaction behaves as an antiferromagnetically ordered object and changes its sign when one moves from one magnetic sublattice to another. Nevertheless, this sign change of the SO interaction can be compensated in some local coordinate frame.
@@
Indeed, the $\{\mathcal{S}|{\bf t}\}$ symmetry justifies the use of the generalized Bloch theorem, which is widely employed in calculations of incommensurate spin spirals without the SO coupling [28, 49]. Using generalized Bloch theorem, any such spiral with the propagation vector $\boldsymbol{q}$ can be described within the crystallographic unit cell. The AFM alignment of spins is a particular case of the spin spiral. However, the generalized Bloch theorem is typically at odds with the relativistic SO interaction, which tends to lock the spins along the easy axis and, thus, severely restrict their ability to rotate in the spin spiral [64]. The antiferromagnetism appears to be a special case: even though the generalized Bloch theorem is not valid for an arbitrary propagation $\boldsymbol{q}$, it can be valid for the AFM one, where the antiferroelectric lattice distortion imposes a special symmetry constrain on the form of the SO interaction. As the result, in the local coordinate frame, the SO interaction and Néel field become translationally invariant on the “ferromagnetic” lattice with one magnetic site in cell.
@@
The regular AFM order doubles the unit cell. In unconventional antiferromagnets with broken $\mathcal{T}$, the magnetic unit cell is expected to coincide with the crystallographic one [1]. We substantially revise this canonical point of view and show that, due to the $\{\mathcal{S}|{\bf t}\}$ symmetry, the magnetic unit cell of centrosymmetric antiferromagnets can be even smaller than the crystallographic one. The situation is highly unusual, but reflects the fundamental symmetry of these systems and naturally explains their similarity with ferromagnets. The emergence of AHE and orbital magnetization in this case becomes natural and they can be evaluated in the same way as for the regular ferromagnets, but in the local coordinate frame, specified by the $\{\mathcal{S}|{\bf t}\}$ symmetry.
@@
The $\{\mathcal{S}|{\bf t}\}$ symmetry enforces the degeneracy of AFM bands and, therefore, has nothing to do with the conventional altermagnetism. From the viewpoint of symmetry, these are different phenomena: the $\{\mathcal{S}|{\bf t}\}$ symmetry is a consequence of inversional invariance, while the band splitting occurs due to rotational $\{\mathcal{C}|{\bf t}\}$ symmetry, which is combined with $\mathcal{T}$. $\{\mathcal{C}|{\bf t}\}$ does not conflict with the $\{\mathcal{S}|{\bf t}\}$ symmetry of the SO interaction, which operates between sites belonging to different sublattices. The main effect of $\{\mathcal{C}|{\bf t}\}$ on the SO interaction is to fix the signs of different components of this interaction around each magnetic site. However, $\{\mathcal{C}|{\bf t}\}$ also specifies the form of hopping parameters operating within the sublattices by transforming the hoppings in one magnetic sublattice to “rotated” hoppings in another sublattice. This can split the AFM bands and eventually break the $\{\mathcal{S}|{\bf t}\}$ symmetry.
@@
If microscopic Hamiltonian were real, the transformation $\{\mathcal{S}|{\bf t}\}$ would be identical to $\{\mathcal{T}|{\bf t}\}$. However, when the bonds connecting different sublattices are noncentrosymmetric, as in the antiferroelectrically distorted lattice, the Hamiltonian is complex. Then, $\{\mathcal{S}|{\bf t}\}$ is not the same as $\{\mathcal{T}|{\bf t}\}$. Therefore, the time-reversal symmetry can be broken, but the Hamiltonian remain invariant under $\{\mathcal{S}|{\bf t}\}$, making the bands degenerate. This is the fundamental reason why the centrosymmetric antiferromagnets can exhibit AHE and other ferromagnetic phenomena.
The basic idea was illustrated on a number of examples, including square perovskite lattice, monoclinic VF4 and CuF2, and antiferromagnetism in the rutile structure. However, it has much wider implication to the properties of unconventional antiferromagnets. The authors of recent review article [23] have proposed 221 three-dimensional materials, which could potentially be altermagnetic, and this number continues to grow. However, most of these materials are centrosymmetric antiferromagnets, whose fundamental ferromagnetic properties are consequences of the hidden $\{\mathcal{S}|{\bf t}\}$ symmetry.
The band splitting provides an additional contribution to AHE. However, in model considerations, it emerges only as a small correction to AHE obtained for the spin-degenerate bands [26]. Moreover, for certain materials, like La2CuO4, the band splitting vanishes due to the special symmetry constraint imposed on parameters of the one-orbital model, which is expected to provide an adequate description near the Fermi level [67, 68].
In the light of these arguments, the systematic experimental information establishing a correlation between the band splitting and AHE is highly welcomed. Today, this information is very sparse. For instance, AHE has been experimentally measured in MnTe [71] and attributed to the large band splitting [20, 21, 71]. AHE has been also reported in AFM FeS [72], but no information about the band splitting is currently available. As an opposite example, no indications of the magnon band splitting has been reported in MnF2 [73], which is yet another altermagnetic candidate crystallizing in the rutile structure [23]. Although this is not a direct probe of the band splitting, some relationship is anticipated because interatomic exchange interactions should reflect details of the electronic structure [30]. Then, an interesting question is whether MnF2 should exhibit the magneto-optical effect or AHE under doping. Our work clearly shows that it should and the ideas of the toy model analysis can be applied in realistic electronic structure calculations.
@@
Thus, the altermagnetic splitting of bands is certainly an interesting theoretical discovery leading to a number of practically important phenomena such as the spin-current generation in AFM substances [15, 18, 24]. However, regarding the whole spectrum of properties expected in centrosymmetric antiferromagnets with broken time-reversal symmetry, it would not be right to attribute all of them to the band splitting and try to consider all of them only from the viewpoint of this splitting. Even if the band splitting is small, the material is still expected to host robust AHE and net orbital magnetization, exhibit magneto-optical rotations and other “ferromagnetic” phenomena, which are related to other, hidden, symmetries of such AFM state, being the consequence of the inversional invariance.